# More on soundness in the enriched context

### v2

@@ 3.2. Characterization theorems
${\mathcal{K}}:=\Phi\textnormal{-}\mathbf{Cont}({\mathcal{C}},{\mathcal{V}})\simeq\Phi\textnormal{-}\mathbf{Cont}({\mathcal{D}},{\mathcal{V}})$
is locally $\Phi$-presentable by Theorem 3.9. By the arguments above, is then enough to prove that ${\mathcal{K}}_{\Phi}$ consists just of the representable functors, that is: ${\mathcal{K}}_{\Phi}\simeq{\mathcal{C}}^{\textnormal{op}}$. On one hand every representable $YC={\mathcal{C}}(C,-)$ in ${\mathcal{K}}$ is clearly $\Phi$-presentable since ${\mathcal{K}}(YC,-)\cong\textnormal{ev}_{C}\circ J$, where $J\colon\Phi\textnormal{-}\mathbf{Cont}({\mathcal{C}},{\mathcal{V}})\to[{\mathcal{C}},{\mathcal{V}}]$ is the inclusion, and both $\textnormal{ev}_{C}$ and $J$ preserves $\Phi$-flat colimits. On the other hand, let $X\in{\mathcal{K}}_{\Phi}$; then $[{\mathcal{C}},{\mathcal{V}}](JX,J-)$ preserves $\Phi$-flat colimits. By [27, Proposition 6.14], $JX$ is representable if and only if $[{\mathcal{C}},{\mathcal{V}}](JX,-)$ preserves the colimit $JX*Y$ (since ${\mathcal{C}}$ is Cauchy complete). But $JX$ is $\Phi$-continuous, and hence $\Phi$-flat; thus $JX*Y\cong J(X*H)$ where $H$ is the inclusion of ${\mathcal{C}}$ in ${\mathcal{K}}$. Thus the thesis follows from the fact that $[{\mathcal{C}},{\mathcal{V}}](JX,J-)$ preserves $\Phi$-flat colimits. ∎
@@
The following theorem then characterizes locally $\Phi$-presentable ${\mathcal{V}}$-categories. For one implication we need the further assumption of ${\mathcal{V}}_{0}$ being locally bounded [24, Section 6.1]. We don’t really need to recall the definition here; it will suffice to say that locally bounded categories include all locally presentable categories as well as certain categories of topological spaces (see [34] for a good amount of examples).
@@
$(3)\Rightarrow(1)$. Assume that ${\mathcal{V}}_{0}$ is locally bounded and let ${\mathcal{M}}$ be a family of morphisms which defines ${\mathcal{K}}={\mathcal{M}}^{\perp}$ as a $\Phi$-orthogonality class in $[{\mathcal{C}},{\mathcal{V}}]$. It follows at once that ${\mathcal{K}}$ is closed in $[{\mathcal{C}},{\mathcal{V}}]$ under $\Phi$-flat colimits and all limits. Then by [24, Theorem 6.5] (since ${\mathcal{V}}_{0}$ is assumed to be locally bounded) the inclusion $J\colon{\mathcal{K}}\hookrightarrow[{\mathcal{C}},{\mathcal{V}}]$ has a left adjoint $F$. It follows that ${\mathcal{K}}$ is locally $\Phi$-presentable by Theorem 3.9. ∎
@@
$[{\mathcal{E}},J]\colon[{\mathcal{E}},{\mathcal{K}}]\hookrightarrow[{\mathcal{E}},[{\mathcal{C}},{\mathcal{V}}]]\cong[{\mathcal{E}}\otimes{\mathcal{C}},{\mathcal{V}}]$
@@
(3). Consider $F_{1}\colon{\mathcal{K}}_{1}\to{\mathcal{L}}$ and $F_{2}\colon{\mathcal{K}}_{2}\to{\mathcal{L}}$ in $\mathbf{L}\Phi\mathbf{P}\to{\mathcal{V}}\textnormal{-}\mathbf{CAT}$ where $F_{2}$ is an isofibration. Consider their pullback ${\mathcal{M}}$ together with the projection $G_{i}\colon{\mathcal{M}}\to{\mathcal{K}}_{1}$. By [32, Theorem 5.5] the ${\mathcal{V}}$-category ${\mathcal{M}}$ is accessible; moreover, since the pullback is also a bipullback ($F_{2}$ is an isofibration) and $F_{1}$ and $F_{2}$ are continuous and preserve $\Phi$-flat colimits; then ${\mathcal{M}}$ is complete and $\Phi$-cocomplete and $G_{1}$ and $G_{2}$ preserve such limits and colimits. It follows that ${\mathcal{M}}$ is locally presentable and $G_{1}\times G_{2}\colon{\mathcal{M}}\to{\mathcal{K}}_{1}\times{\mathcal{K}}_{2}$ is conservative, preserves $\Phi$-flat colimits, and is a right adjoint (being continuous between locally presentable ${\mathcal{V}}$-categories). Thus ${\mathcal{M}}$ is locally $\Phi$-presentable by Proposition 3.4. This implies that ${\mathcal{M}}$ is also the pullback of $F_{1}$ and $F_{2}$ in $\mathbf{L}\Phi\mathbf{P}\to{\mathcal{V}}\textnormal{-}\mathbf{CAT}$. ∎
@@
Follows directly from Lemma 3.19 above since products, powers, and pullbacks along isofibrations, together with splittings of idempotent equivalences (which exist by [11, Remark 7.6]), imply the existence of all flexible limits (argue as in the beginning of [32, Theorem 5.5]). ∎
@@ 3.3. Sound classes induced from $\mathbf{Set}$
([29, Section 5]) Denote by $\Phi_{\mathbb{D}}$ the locally small class of weights in ${\mathcal{V}}$ generated by the conical weights in ${\mathbb{D}}$ and powers by objects in $({\mathcal{V}}_{0})_{\mathbb{D}}$.
@@
([29, 5.22]). Let ${\mathcal{C}}$ be small and $\Phi_{\mathbb{D}}$-cocomplete. The following are equivalent for a weight $M\colon{\mathcal{C}}^{\textnormal{op}}\to{\mathcal{V}}$:
@@
The proof is exactly the same as that of [32, Proposition 3.20], simply replace $\alpha$-small colimits with $\Phi_{\mathbb{D}}$-colimits and use Proposition 3.25 above. ∎
@@
Since the class of $\alpha$-small weights is pre-saturated [42, Definition A.1.13], this means that every Cauchy weight $M$ is the left Kan extension of some $\alpha$-small weight [42, Proposition A.1.14].
@@
If ${\mathcal{V}}_{0}$ is locally presentable then it is also locally $\alpha$-presentable as a closed category for some $\alpha$ by [26, Proposition 2.4]. Thus the result follows from the corollary above since the class of $\alpha$-small weights is essentially small. ∎
@@
The corollary above applies when ${\mathcal{V}}$ is locally dualizable category as in [31, Definition 4.3]. Indeed, one needs to take ${\mathbb{D}}$ to be the class of finite products; then $\Phi_{\mathbb{D}}$-colimits are generated by finite direct sums and powers by a strong generator made of dualizable objects. Thus, as a consequence of the Corollary above we recover [31, Corollary 4.22], and in particular the characterization of absolutely complete $\operatorname{\mathbf{Ab}}$-categories and $\mathbf{GAb}$-categories ([37]).
@@ 4. Soundness and raising the index of accessibility
Unlike the locally presentable case, raising the index of accessibility is an issue: ordinarily it is not true that if ${\mathcal{A}}$ is an $\alpha$-accessible category and $\beta>\alpha$, then ${\mathcal{A}}$ is $\beta$-accessible. This problem was address from the very beginning [36, 4] by introducing a new order relationship $\triangleleft$, called the sharply less relationship, between regular cardinals. In [36, Section 2.3] it is proved that, given $\alpha<\beta$, every $\alpha$-accessible category ${\mathcal{A}}$ is $\beta$-accessible if and only if $\alpha\triangleleft\beta$, and in that case every $\beta$-presentable object of ${\mathcal{A}}$ is a $\beta$-small $\alpha$-filtered colimit of $\alpha$-presentable objects.
In the context of a sound class of weights this problem has never been addressed (not even for ordinary categories); all that is known is that every $\Phi$-accessible ${\mathcal{V}}$-category ${\mathcal{A}}$ is $\alpha$-accessible for arbitrarily large regular cardinals $\alpha$ [33, Theorem 4.4], but such $\alpha$ might depend on the given ${\mathcal{V}}$-category ${\mathcal{A}}$. In this section we shall extend the sharply less relationship to the context of sound classes of weights and prove theorems similar to those known in the ordinary setting.
@@
[…]
@@
coincides with the Cauchy completion of ${\mathcal{C}}$. This follows from a straightforward generalization of [41, Lemma 3.3] where instead of $\alpha$-small weights one uses the sound class $\Phi$.
@@
as full subcategories of ${\mathcal{P}}{\mathcal{C}}$. Indeed, $\Psi\negthinspace^{{}^{+}}\negthinspace(\Phi^{{}^{+}}_{\Psi}{\mathcal{C}})$ is contained in ${\mathcal{P}}{\mathcal{C}}$ by [33, 4.10] since $\Psi\negthinspace^{{}^{+}}\negthinspace$-colimits commute in ${\mathcal{V}}$ with $\Phi^{{}^{+}}_{\Psi}$-limits (these limits are in particular $\Psi$-limits), and then $\Psi\negthinspace^{{}^{+}}\negthinspace(\Phi^{{}^{+}}_{\Psi}{\mathcal{C}})\subseteq\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}$ since both $\Psi\negthinspace^{{}^{+}}\negthinspace$ and $\Phi^{{}^{+}}_{\Psi}$ are contained in $\Phi\negthinspace^{{}^{+}}\negthinspace$.
@@
We are now ready to state and prove the following result which generalizes [33, Theorem 2.3] to general sound classes of weights.
@@
    For any small ${\mathcal{V}}$-category ${\mathcal{C}}$ the inclusion (1) above is an equivalence
    $\Psi\negthinspace^{{}^{+}}\negthinspace(\Phi^{{}^{+}}_{\Psi}{\mathcal{C}})\simeq\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}.$
@@
    Every $\Phi$-accessible ${\mathcal{V}}$-category ${\mathcal{A}}$ is $\Psi$-accessible and the inclusion (2) above is an equivalence
    $\Phi^{{}^{+}}_{\Psi}{\mathcal{A}}_{\Phi}\simeq{\mathcal{A}}_{\Psi}.$
@@
Now, by the proof of [32, 3.9] it follows that ${\mathcal{A}}_{\Psi}$ is the Cauchy completion of $\Phi^{{}^{+}}_{\Psi}{\mathcal{A}}_{\Phi}$, which is already Cauchy complete. Thus ${\mathcal{A}}_{\Psi}\simeq\Phi^{{}^{+}}_{\Psi}{\mathcal{A}}_{\Phi}$.
@@
When $\Phi$ and $\Psi$ are respectively ${\mathcal{P}}_{\alpha}$ and ${\mathcal{P}}_{\beta}$, for $\alpha<\beta$, then it is enough to require that $\alpha$-accessibility implies $\beta$-accessibility to obtain that ${\mathcal{P}}_{\alpha}\triangleleft{\mathcal{P}}_{\beta}$: the additional conditions in (3) and (4) are then automatically satisfied (see [36, Proposition 2.3.11] for the ordinary result; the enriched version follows from [32, Theorem 3.25]). It follows that ${\mathcal{P}}_{\alpha}\triangleleft{\mathcal{P}}_{\beta}$ if and only if $\alpha\triangleleft\beta$ as regular cardinals.
@@
For that, consider the first $\beta^{\prime}$ such that $\Phi\subseteq{\mathcal{P}}_{\beta^{\prime}}$. Then, there exists $\beta\triangleright\beta^{\prime}$ for which ${\mathcal{P}}_{\beta^{\prime}}{\mathcal{C}}\subseteq\Phi^{{}^{+}}_{\beta}(\Phi{\mathcal{C}})$ for any ${\mathcal{C}}$. Indeed, fix a small set $S$ of weights whose saturation is ${\mathcal{P}}_{\beta^{\prime}}$, then since $\Phi\negthinspace^{{}^{+}}\negthinspace(\Phi(-))\simeq{\mathcal{P}}(-)$ (by [33, Proposition 4.9]), it follows that any weight $M\in S$ can be written as a $\Phi$-flat colimit of $\Phi$-colimits of representables. The $\Phi$-flat colimit involved in such description of $M$ will be $\beta_{M}$-small for sone $\beta_{M}$; thus it is enough to consider $\beta:=\textnormal{sup}_{M\in S}\beta_{M}$. It is now clear that for such $\beta$ we have ${\mathcal{P}}_{\beta^{\prime}}{\mathcal{C}}\subseteq\Phi^{{}^{+}}_{\beta}(\Phi{\mathcal{C}})$ for any ${\mathcal{C}}$.
@@
holds for any ${\mathcal{C}}$. For the converse note that, since $\beta^{\prime}\triangleright\beta$, all $\beta$-small conical colimits are generated by the $\beta^{\prime}$-small and the $\beta$-small $\beta^{\prime}$-filtered conical ones. It follows that $\beta$-small copowers can also be written as $\beta$-small $\beta^{\prime}$-filtered conical colimits of $\beta^{\prime}$-small copowers. Therefore, any object of ${\mathcal{P}}_{\beta}{\mathcal{C}}$ (that is, any $\beta$-small weighted colimit of representables) can be rewritten as a $\beta$-small $\beta^{\prime}$-filtered colimit from ${\mathcal{P}}_{\beta^{\prime}}{\mathcal{C}}$. But ${\mathcal{P}}_{\beta^{\prime}}{\mathcal{C}}\subseteq\Phi^{{}^{+}}_{\beta}(\Phi{\mathcal{C}})$ and $\beta$-small $\beta^{\prime}$-filtered colimits are in $\Phi^{{}^{+}}_{\beta}$ (since $\Phi\subseteq{\mathcal{P}}_{\beta^{\prime}}$). Thus $\Phi^{{}^{+}}_{\beta}(\Phi{\mathcal{C}})={\mathcal{P}}_{\beta}{\mathcal{C}}$.
@@
The argument is similar to that of [32, 3.25]. Let ${\mathcal{C}}$ be any small ${\mathcal{V}}$-category, then we can see $\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}$ as the intersection
@@
where $H\colon{\mathcal{C}}\hookrightarrow\Phi{\mathcal{C}}$ is the inclusion. Indeed, a ${\mathcal{V}}$-functor $M\colon{\mathcal{C}}^{\textnormal{op}}\to{\mathcal{V}}$ lies in $\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}$ if and only if $M$ is $\Phi$-flat, if and only if $\textnormal{Lan}_{H^{\textnormal{op}}}M$ is $\Phi$-flat (by [32, 3.3]), if and only if $\textnormal{Lan}_{H^{\textnormal{op}}}M$ is $\Phi$-continuous (by soundness of $\Phi$), if and only if $\textnormal{Lan}_{H^{\textnormal{op}}}M\cong\textnormal{Ran}_{H^{\textnormal{op}}}M$ (since $(\Phi{\mathcal{C}})^{\textnormal{op}}$ is the free completion of ${\mathcal{C}}^{\textnormal{op}}$ under $\Phi$-colimits).
By [32, Lemma 3.24] and its proof, to show that $\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}$ is $\alpha$-accessible and that $J$ preserves the $\alpha$-presentable objects, it is enough to show that $\textnormal{Ran}_{H^{\textnormal{op}}}$ and $\textnormal{Lan}_{H^{\textnormal{op}}}$ both preserve $\beta$-flat colimits and the $\beta$-presentable objects.
@@
    If $\Phi={\mathcal{P}}_{\lambda}$ we can choose $\beta=\lambda$; hence we obtain that ${\mathcal{P}}_{\lambda}\triangleleft{\mathcal{P}}_{\alpha}$ whenever $\lambda\triangleleft\alpha$. (This is also a consequence of [32, 3.25].)
@@
This follows from the theorem above, the fact that $\triangleleft$ is transitive, and that, given $\alpha$, there are arbitrarily large $\lambda$ with $\alpha\triangleleft\lambda$ (see [36, Corollary 2.3.6]). ∎

### v3

@@ 3.2. Characterization theorems
${\mathcal{K}}:=\Phi\textnormal{-}\textnormal{Cont}({\mathcal{C}},{\mathcal{V}})\simeq\Phi\textnormal{-}\textnormal{Cont}({\mathcal{D}},{\mathcal{V}})$
is locally $\Phi$-presentable by Theorem 3.9. By the arguments above, is then enough to prove that ${\mathcal{K}}_{\Phi}$ consists just of the representable functors, that is: ${\mathcal{K}}_{\Phi}\simeq{\mathcal{C}}^{\textnormal{op}}$. On one hand every representable $YC={\mathcal{C}}(C,-)$ in ${\mathcal{K}}$ is clearly $\Phi$-presentable since ${\mathcal{K}}(YC,-)\cong\textnormal{ev}_{C}\circ J$, where $J\colon\Phi\textnormal{-}\textnormal{Cont}({\mathcal{C}},{\mathcal{V}})\to[{\mathcal{C}},{\mathcal{V}}]$ is the inclusion, and both $\textnormal{ev}_{C}$ and $J$ preserves $\Phi$-flat colimits. On the other hand, let $X\in{\mathcal{K}}_{\Phi}$; then $[{\mathcal{C}},{\mathcal{V}}](JX,J-)$ preserves $\Phi$-flat colimits. By [28, Proposition 6.14], $JX$ is representable if and only if $[{\mathcal{C}},{\mathcal{V}}](JX,-)$ preserves the colimit $JX*Y$ (since ${\mathcal{C}}$ is Cauchy complete). But $JX$ is $\Phi$-continuous, and hence $\Phi$-flat; thus $JX*Y\cong J(X*H)$ where $H$ is the inclusion of ${\mathcal{C}}$ in ${\mathcal{K}}$. Thus the thesis follows from the fact that $[{\mathcal{C}},{\mathcal{V}}](JX,J-)$ preserves $\Phi$-flat colimits. ∎
@@
The following theorem then characterizes locally $\Phi$-presentable ${\mathcal{V}}$-categories. For one implication we need the further assumption of ${\mathcal{V}}_{0}$ being locally bounded [25, Section 6.1]. We don’t really need to recall the definition here; it will suffice to say that locally bounded categories include all locally presentable categories as well as certain categories of topological spaces (see [35] for a good amount of examples).
@@
$(3)\Rightarrow(1)$. Assume that ${\mathcal{V}}_{0}$ is locally bounded and let ${\mathcal{M}}$ be a family of morphisms which defines ${\mathcal{K}}={\mathcal{M}}^{\perp}$ as a $\Phi$-orthogonality class in $[{\mathcal{C}},{\mathcal{V}}]$. It follows at once that ${\mathcal{K}}$ is closed in $[{\mathcal{C}},{\mathcal{V}}]$ under $\Phi$-flat colimits and all limits. Then by [25, Theorem 6.5] (since ${\mathcal{V}}_{0}$ is assumed to be locally bounded) the inclusion $J\colon{\mathcal{K}}\hookrightarrow[{\mathcal{C}},{\mathcal{V}}]$ has a left adjoint $F$. It follows that ${\mathcal{K}}$ is locally $\Phi$-presentable by Theorem 3.9. ∎
@@
${\mathcal{E}},J]\colon[{\mathcal{E}},{\mathcal{K}}]\hookrightarrow[{\mathcal{E}},[{\mathcal{C}},{\mathcal{V}}]]\cong[{\mathcal{E}}\otimes{\mathcal{C}},{\mathcal{V}}]$
@@
(3). Consider $F_{1}\colon{\mathcal{K}}_{1}\to{\mathcal{L}}$ and $F_{2}\colon{\mathcal{K}}_{2}\to{\mathcal{L}}$ in $\mathbf{L}\Phi\mathbf{P}\to{\mathcal{V}}\textnormal{-}\mathbf{CAT}$ where $F_{2}$ is an isofibration. Consider their pullback ${\mathcal{M}}$ together with the projection $G_{i}\colon{\mathcal{M}}\to{\mathcal{K}}_{1}$. By [33, Theorem 5.5] the ${\mathcal{V}}$-category ${\mathcal{M}}$ is accessible; moreover, since the pullback is also a bipullback ($F_{2}$ is an isofibration) and $F_{1}$ and $F_{2}$ are continuous and preserve $\Phi$-flat colimits; then ${\mathcal{M}}$ is complete and $\Phi$-cocomplete and $G_{1}$ and $G_{2}$ preserve such limits and colimits. It follows that ${\mathcal{M}}$ is locally presentable and $G_{1}\times G_{2}\colon{\mathcal{M}}\to{\mathcal{K}}_{1}\times{\mathcal{K}}_{2}$ is conservative, preserves $\Phi$-flat colimits, and is a right adjoint (being continuous between locally presentable ${\mathcal{V}}$-categories). Thus ${\mathcal{M}}$ is locally $\Phi$-presentable by Proposition 3.4. This implies that ${\mathcal{M}}$ is also the pullback of $F_{1}$ and $F_{2}$ in $\mathbf{L}\Phi\mathbf{P}\to{\mathcal{V}}\textnormal{-}\mathbf{CAT}$. ∎
@@
Follows directly from Lemma 3.21 above since products, powers, and pullbacks along isofibrations, together with splittings of idempotent equivalences (which exist by [11, Remark 7.6]), imply the existence of all flexible limits (argue as in the beginning of [33, Theorem 5.5]). ∎
@@ 3.3. Sound classes induced from $\mathbf{Set}$
([30, Section 5]) Denote by $\Phi_{\mathbb{D}}$ the locally small class of weights in ${\mathcal{V}}$ generated by the conical weights in ${\mathbb{D}}$ and powers by objects in $({\mathcal{V}}_{0})_{\mathbb{D}}$.
@@
([30, 5.22]). Let ${\mathcal{C}}$ be small and $\Phi_{\mathbb{D}}$-cocomplete. The following are equivalent for a weight $M\colon{\mathcal{C}}^{\textnormal{op}}\to{\mathcal{V}}$:
@@
The proof is exactly the same as that of [33, Proposition 3.20], simply replace $\alpha$-small colimits with $\Phi_{\mathbb{D}}$-colimits and use Proposition 3.27 above. ∎
@@
Since the class of $\alpha$-small weights is pre-saturated [44, Definition A.1.13], this means that every Cauchy weight $M$ is the left Kan extension of some $\alpha$-small weight [44, Proposition A.1.14].
@@
If ${\mathcal{V}}_{0}$ is locally presentable then it is also locally $\alpha$-presentable as a closed category for some $\alpha$ by [27, Proposition 2.4]. Thus the result follows from the corollary above since the class of $\alpha$-small weights is essentially small. ∎
@@
The corollary above applies when ${\mathcal{V}}$ is locally dualizable category as in [32, Definition 4.3]. Indeed, one needs to take ${\mathbb{D}}$ to be the class of finite products; then $\Phi_{\mathbb{D}}$-colimits are generated by finite direct sums and powers by a strong generator made of dualizable objects. Thus, as a consequence of the Corollary above we recover [32, Corollary 4.22], and in particular the characterization of absolutely complete $\operatorname{\mathbf{Ab}}$-categories and $\mathbf{GAb}$-categories ([39]).
@@ 4. Soundness and raising the index of accessibility
Unlike the locally presentable case, raising the index of accessibility is an issue: ordinarily it is not true that if ${\mathcal{A}}$ is an $\alpha$-accessible category and $\beta>\alpha$, then ${\mathcal{A}}$ is $\beta$-accessible. This problem was address from the very beginning [38, 4] by introducing a new order relationship $\triangleleft$, called the sharply less relationship, between regular cardinals. In [38, Section 2.3] it is proved that, given $\alpha<\beta$, every $\alpha$-accessible category ${\mathcal{A}}$ is $\beta$-accessible if and only if $\alpha\triangleleft\beta$, and in that case every $\beta$-presentable object of ${\mathcal{A}}$ is a $\beta$-small $\alpha$-filtered colimit of $\alpha$-presentable objects.
In the context of a sound class of weights this problem has never been addressed (not even for ordinary categories); all that is known is that every $\Phi$-accessible ${\mathcal{V}}$-category ${\mathcal{A}}$ is $\alpha$-accessible for arbitrarily large regular cardinals $\alpha$ [34, Theorem 4.4], but such $\alpha$ might depend on the given ${\mathcal{V}}$-category ${\mathcal{A}}$. In this section we shall extend the sharply less relationship to the context of sound classes of weights and prove theorems similar to those known in the ordinary setting.
@@
$\Phi^{{}^{+}}_{\Psi}{\mathcal{C}}\displaystyle\lrcorner$$\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}$$\Psi{\mathcal{C}}$$[{\mathcal{C}}^{\textnormal{op}},{\mathcal{V}}]$
@@
coincides with the Cauchy completion of ${\mathcal{C}}$. This follows from a straightforward generalization of [43, Lemma 3.3] where instead of $\alpha$-small weights one uses the sound class $\Phi$.
@@
as full subcategories of ${\mathcal{P}}{\mathcal{C}}$. Indeed, $\Psi\negthinspace^{{}^{+}}\negthinspace(\Phi^{{}^{+}}_{\Psi}{\mathcal{C}})$ is contained in ${\mathcal{P}}{\mathcal{C}}$ by [34, 4.10] since $\Psi\negthinspace^{{}^{+}}\negthinspace$-colimits commute in ${\mathcal{V}}$ with $\Phi^{{}^{+}}_{\Psi}-$limits (these limits are in particular $\Psi$-limits), and then $\Psi\negthinspace^{{}^{+}}\negthinspace(\Phi^{{}^{+}}_{\Psi}{\mathcal{C}})\subseteq\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}$ since both $\Psi\negthinspace^{{}^{+}}\negthinspace$ and $\Phi^{{}^{+}}_{\Psi}$ are contained in $\Phi\negthinspace^{{}^{+}}\negthinspace$.
@@
We are now ready to state and prove the following result which generalizes [34, Theorem 2.3] to general sound classes of weights.
@@
    For any small ${\mathcal{V}}$-category ${\mathcal{C}}$ the inclusion (1) above is an equivalence Ψ^^+(Φ^^+_ΨC) Φ^^+C.
@@
    Every $\Phi$-accessible ${\mathcal{V}}$-category ${\mathcal{A}}$ is $\Psi$-accessible and the inclusion (2) above is an equivalence Φ^^+_ΨAA_Ψ.
@@
Now, by the proof of [33, 3.9] it follows that ${\mathcal{A}}_{\Psi}$ is the Cauchy completion of $\Phi^{{}^{+}}_{\Psi}{\mathcal{A}}_{\Phi}$, which is already Cauchy complete. Thus ${\mathcal{A}}_{\Psi}\simeq\Phi^{{}^{+}}_{\Psi}{\mathcal{A}}_{\Phi}$.
@@
When $\Phi$ and $\Psi$ are respectively ${\mathcal{P}}_{\alpha}$ and ${\mathcal{P}}_{\beta}$, for $\alpha<\beta$, then it is enough to require that $\alpha$-accessibility implies $\beta$-accessibility to obtain that ${\mathcal{P}}_{\alpha}\triangleleft{\mathcal{P}}_{\beta}$: the additional conditions in (3) and (4) are then automatically satisfied (see [38, Proposition 2.3.11] for the ordinary result; the enriched version follows from [33, Theorem 3.25]). It follows that ${\mathcal{P}}_{\alpha}\triangleleft{\mathcal{P}}_{\beta}$ if and only if $\alpha\triangleleft\beta$ as regular cardinals.
@@
For that, consider the first $\beta^{\prime}$ such that $\Phi\subseteq{\mathcal{P}}_{\beta^{\prime}}$. Then, there exists $\beta\triangleright\beta^{\prime}$ for which ${\mathcal{P}}_{\beta^{\prime}}{\mathcal{C}}\subseteq\Phi^{{}^{+}}_{\beta}(\Phi{\mathcal{C}})$ for any ${\mathcal{C}}$. Indeed, fix a small set $S$ of weights whose saturation is ${\mathcal{P}}_{\beta^{\prime}}$, then since $\Phi\negthinspace^{{}^{+}}\negthinspace(\Phi(-))\simeq{\mathcal{P}}(-)$ (by [34, Proposition 4.9]), it follows that any weight $M\in S$ can be written as a $\Phi$-flat colimit of $\Phi$-colimits of representables. The $\Phi$-flat colimit involved in such description of $M$ will be $\beta_{M}$-small for sone $\beta_{M}$; thus it is enough to consider $\beta\geq\textnormal{sup}_{M\in S}\beta_{M}$. It is now clear that for such $\beta$ we have ${\mathcal{P}}_{\beta^{\prime}}{\mathcal{C}}\subseteq\Phi^{{}^{+}}_{\beta}(\Phi{\mathcal{C}})$ for any ${\mathcal{C}}$.
@@
holds for any ${\mathcal{C}}$. For the converse note that, since $\beta\triangleright\beta^{\prime}$, all $\beta$-small conical colimits are generated by the $\beta^{\prime}$-small and the $\beta$-small $\beta^{\prime}$-filtered conical ones. It follows that $\beta$-small copowers can also be written as $\beta$-small $\beta^{\prime}$-filtered conical colimits of $\beta^{\prime}$-small copowers. Therefore, any object of ${\mathcal{P}}_{\beta}{\mathcal{C}}$ (that is, any $\beta$-small weighted colimit of representables) can be rewritten as a $\beta$-small $\beta^{\prime}$-filtered colimit from ${\mathcal{P}}_{\beta^{\prime}}{\mathcal{C}}$. But ${\mathcal{P}}_{\beta^{\prime}}{\mathcal{C}}\subseteq\Phi^{{}^{+}}_{\beta}(\Phi{\mathcal{C}})$ and $\beta$-small $\beta^{\prime}$-filtered colimits are in $\Phi^{{}^{+}}_{\beta}$ (since $\Phi\subseteq{\mathcal{P}}_{\beta^{\prime}}$). Thus $\Phi^{{}^{+}}_{\beta}(\Phi{\mathcal{C}})={\mathcal{P}}_{\beta}{\mathcal{C}}$.
@@
The argument is similar to that of [33, 3.25]. Let ${\mathcal{C}}$ be any small ${\mathcal{V}}$-category, then we can see $\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}$ as the intersection
@@
where $H\colon{\mathcal{C}}\hookrightarrow\Phi{\mathcal{C}}$ is the inclusion. Indeed, a ${\mathcal{V}}$-functor $M\colon{\mathcal{C}}^{\textnormal{op}}\to{\mathcal{V}}$ lies in $\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}$ if and only if $M$ is $\Phi$-flat, if and only if $\textnormal{Lan}_{H^{\textnormal{op}}}M$ is $\Phi$-flat (by [33, 3.3]), if and only if $\textnormal{Lan}_{H^{\textnormal{op}}}M$ is $\Phi$-continuous (by soundness of $\Phi$), if and only if $\textnormal{Lan}_{H^{\textnormal{op}}}M\cong\textnormal{Ran}_{H^{\textnormal{op}}}M$ (since $(\Phi{\mathcal{C}})^{\textnormal{op}}$ is the free completion of ${\mathcal{C}}^{\textnormal{op}}$ under $\Phi$-colimits).
By [33, Lemma 3.24] and its proof, to show that $\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}$ is $\alpha$-accessible and that $J$ preserves the $\alpha$-presentable objects, it is enough to show that $\textnormal{Ran}_{H^{\textnormal{op}}}$ and $\textnormal{Lan}_{H^{\textnormal{op}}}$ both preserve $\beta$-flat colimits and the $\beta$-presentable objects.
@@
    If $\Phi={\mathcal{P}}_{\lambda}$ we can choose $\beta=\lambda$; hence we obtain that ${\mathcal{P}}_{\lambda}\triangleleft{\mathcal{P}}_{\alpha}$ whenever $\lambda\triangleleft\alpha$. (This is also a consequence of [33, 3.25].)
@@
This follows from the theorem above, the fact that $\triangleleft$ is transitive, and that, given $\alpha$, there are arbitrarily large $\lambda$ with $\alpha\triangleleft\lambda$ (see [38, Corollary 2.3.6]). ∎